OpenStudy (anonymous):

If one root of the equation x^2+kx-8=0 is the square of the other, what is k? A.) -6 B.) -2 C.) 2 D.) 6 E.) 8

OpenStudy (alivejeremy):

x=k ± − k2+32−−−−−−√ 2

hartnn (hartnn):

so can we take a, a^2 as the 2 roots ?

OpenStudy (alivejeremy):

So the answer would be C

OpenStudy (alivejeremy):

@erika_riera

OpenStudy (anonymous):

I put the answer c the first time but I got it wrong.

OpenStudy (alivejeremy):

Well try B

OpenStudy (alivejeremy):

Hold on

hartnn (hartnn):

why are you ppl guessing instead of solving? :P

OpenStudy (alivejeremy):

Yeah thats why i said hold on ^^^:P

OpenStudy (anonymous):

I'm doing test corrections and I need someone to explain to me how to solve it

hartnn (hartnn):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hartnn so can we take a, a^2 as the 2 roots ? \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (anonymous):

yes

hartnn (hartnn):

product of roots for \(ax^2+bx+c = 0\) is = c/a so product of roots in your case = a*a^2 = -8/1 does this make sense?

OpenStudy (anonymous):

yes I get that

hartnn (hartnn):

a^3 = -8 find a

OpenStudy (anonymous):

where did you get the ^3 from?

hartnn (hartnn):

\(\Large a \times a^2 = a^{1+2} = a^3\)

OpenStudy (anonymous):

a=2?

hartnn (hartnn):

2^3 = 8 not -8

OpenStudy (anonymous):

a=-2?

hartnn (hartnn):

yes!

hartnn (hartnn):

so a = -2 is one of the root now you can just plug in x= -2 in your original equation to get k

hartnn (hartnn):

x^2+kx-8=0 (-2)^2 -2k - 8 =0 get k

OpenStudy (anonymous):

So k=-2?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

oh okay I get it now thank you!

hartnn (hartnn):

welcome ^_^